OpenStudy (anonymous):

true or false sinx cosy = 1/2(sin(x+y)+ cos(x-y)) answer with explaination please

OpenStudy (anonymous):

sin(x+y) = sin(x)cos(y) + sin(y)cos(x) cos(x-y) = cos(x)cos(y) + sin(x)sin(y) can you answer now ?

OpenStudy (anonymous):

its true becuase 1/2 of 2 sinxcosy is that?

OpenStudy (anonymous):

show me .. use the formulas that i gave and show me that the right side equal the left side if it's true you should be able to show it

OpenStudy (anonymous):

er ..i kinda got it in my head ..i just can't put in on paper ...can you help me ._.

OpenStudy (anonymous):

sinx cosy = 1/2(sin(x)cos(y) + sin(y)cos(x)+ cos(x)cos(y) + sin(x)sin(y))

OpenStudy (anonymous):

try working with this now

OpenStudy (anonymous):

her ..multiply by 2 on both sides?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

try manipulating it .. do you think it is true ?

OpenStudy (anonymous):

are you there ?

OpenStudy (anonymous):

yeah sorry im uh trying ..but confused alittle

OpenStudy (anonymous):

if you cant show that the right side is equal to the left side after manipulating it might be false! in fact we can show that this is false by choosing x=0 and y=0 and put it in the original equation : sinx cosy = 1/2(sin(x+y)+ cos(x-y)) sin0cos0 = 1/2(sin0 + cos0) 0 = 1/2(0 + 1) 0 = 1/2 false!

OpenStudy (anonymous):

oh ._. i got it backwards

OpenStudy (anonymous):

note that in order to show that this is true we cant just choose numbers we really have to manipulate the original equation and see that the left side equals the right side but if we want to show that it is false choosing two values is enough!

OpenStudy (anonymous):

oh okay ty